# Comparison results for Markov tree distributions

### v1

@@ Abstract
We develop comparison results for Markov tree distributions extending ordering results from the literature on discrete time Markov processes and recently studied ordering results for conditionally independent factor models to tree structures. Based on fairly natural positive dependence conditions, our main contribution is a comparison result with respect to the supermodular order. Since this order is a pure dependence order, it has many applications in optimal transport, finance, and insurance. As an illustrative example, we consider hidden Markov models and study distributional robustness for functionals of the random walk under model uncertainty. Further, we show that, surprisingly, more general comparison results via the recently established rearrangement-based Schur order for conditional distributions, which implies an ordering of Chatterjee’s rank correlation, do not carry over from star structures to trees. Several examples and a detailed discussion of the assumptions demonstrate the generality of our results and provide further insights into the behavior of multidimensional distributions.
Keywords convex risk measure, distributional robustness, factor model, hidden Markov model, Markov process, optimal transport, positive dependence, supermodular order, stochastically increasing, vine copula model.
@@ 1 Introduction
Let $T=(N,E)$ be a tree with finitely or countably many nodes $N=\{0,1,2,\ldots\}$ and edges $E\subset N\times N\,.$ A sequence $(X_{n})_{n\in N}$ of random variables is said to follow a Markov tree distribution with respect to $T$ if, for each two finite and disjoint sets $A,B\subset N\,,$ the variables $(X_{a})_{a\in A}$ and $(X_{b})_{b\in B}$ are conditionally independent given $X_{i}$ for any node $i\in N$ that separates $A$ and $B,$ see Definition 2.3. This concept extends the Markov property from a chain of nodes to tree structures, noting that for Markov processes future and past events are conditionally independent given the present value. Further subclasses are conditionally independent factor models, hidden Markov models [29, 43], tree-indexed Markov chains [22, 89], and vine copula models truncated after the first level [17, 26], see Figure 1. Every Markov tree distribution can be specified by a sequence $F=(F_{n})_{n\in N}$ of univariate marginal distribution functions and a sequence $B=(B_{e})_{e\in E}$ of bivariate copulas that describe the dependence structure between each two random variables that are adjacent in $T\,,$ see Proposition 2.5. We write $\mathcal{M}(F,T,B)$ for the Markov tree distribution with these specifications.
The main contribution of our paper, Theorem 1.3, is a supermodular ordering result for Markov tree distributed random variables $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}\,,$ where we establish general conditions on the specifications of $X$ and $Y$ to infer integral inequalities of the form
$\mathbb{E}[f(X)]\leq\mathbb{E}[f(Y)]$ (1)
for all supermodular functions $f$ such that the expectations exist. The class of supermodular functions includes many interesting functions such as componentwise minima and maxima, distribution and survival functions, as well as convex functions of the component sum, see Table 1. In particular, for any law-invariant convex risk measure $\Psi\,,$ the functional $\Psi(\sum_{n}X_{n})$ is consistent with the supermodular order, which furthermore has the appealing properties that it is a pure dependence order (i.e. $\mathsf{law}(X_{n})=\mathsf{law}(Y_{n})$ for all $n\in N$) and invariant under increasing transformations of the components [87, 101, 104]. Hence, it allows applications in various fields, where the marginal distributions are assumed to be fixed, such as optimal transport or mathematical finance [18, 19, 20, 50, 92, 115, 121]. Integral stochastic orders are useful in various applications, since many parametric models in statistics exhibit monotonicity properties in their parameters for several classes of functionals [82, 85, 101, 120]. The novelty of our approach is that we establish, for the flexible class of Markov tree distributions, general conditions on the marginal and dependence specifications implying integral inequalities as in (1) for several classes of functions, see Section 3. In particular, our results provide simple conditions for constructing and comparing positive supermodular dependent distributions: As building blocks, we can take any set of univariate distribution functions and any set of bivariate copulas that are conditionally increasing and pointwise larger than the product copula. A supermodular comparison based on bivariate tree specifications is then obtained by a pointwise comparison of the bivariate copulas, see Theorem 3.1.
As a consequence, we obtain distributional robustness for various functionals that are consistent with the supermodular order. To this end, consider for a tree $T=(N,E)\,,$ for marginal distribution functions $F=(F_{n})_{n\in N}\,,$ and for suitable subclasses $(\mathcal{C}_{e})_{e\in E}$ of bivariate copulas, the problem
@@
$B$ $=(B_{e})_{e\in E}\text{ with }B_{e}\in\mathcal{C}_{e}\text{ for all }e\in E\,,$ (3)
where $f$ is a supermodular function. Condition (2) ensures that the univariate marginal distributions are fixed and that $X$ satisfies the Markov property with respect to the tree $T$. The copula constraints in (3) relate to the dependence structure of the transition kernels of the Markov tree distributions. Under positive dependence conditions on the classes $\mathcal{C}_{e}$ of bivariate copulas as above, we determine in Corollary 3.2 solutions to the above optimization problem in order to obtain distributional robustness for various functionals. Our comparison results with respect to the lower orthant, upper orthant, and directionally convex order also allow to incorporate distributional robustness in the marginal distributions. In Section 4, we give an application to hidden Markov models, where we determine a dependence uncertainty band for the distribution function of the maximal observations of a random walk under model uncertainty and noise. Note that the above optimization problem can also be interpreted as a multi-marginal Markovian optimal transport problem with dependence constraints. It is related to various optimal transport problems studied in the literature, for instance, to multi-marginal optimal transport, optimal transport with linear constraints, and martingale optimal transport [18, 19, 20, 50, 121]. Our comparison results also allow connections to the literature on similarity of stochastic processes, as studied in the context of adaptive, causal or bicausal optimal transport in [11, 12, 13, 53, 98].
1.1 Main result
[…]
@@ 1.1 Main result
Figure 1: Examples of Markov tree distributions: a) A Markov process in discrete time, where the underlying tree is a chain. b) Random variables $X_{1},\ldots,X_{d}$ that are conditionally independent given the common factor variable $X_{0}\,,$ where the underlying tree has a star-like structure. c) A tree-indexed Markov process with a general underlying tree structure. d) A hidden Markov model with hidden nodes $(X_{i})_{i\in 2\mathbb{N}_{0}}$ and observable nodes $(X_{i})_{i\in 2\mathbb{N}_{0}+1}$. Each model is uniquely determined by the univariate marginal distributions specifying the nodes and by the bivariate copulas $(B_{ij})_{(i,j)\in E}$ specifying the edges of the underlying tree $T=(N,E)\,,$ see Proposition 2.5. If Lebesgue-densities exist, a Markov tree distribution is a vine copula model truncated after the first level, where a) corresponds to a D-vine structure and b) to a C-vine structure, see [26].
@@
While comparison results for Markov processes have been known for a long time [38, 58, 103], a supermodular comparison of star structures has been shown recently in [8]. Both results are based on asymmetric positive dependence conditions, which differ for chain and star structures. Our main comparison result for Markov tree distributions is based on the proof for ordering Markov processes in [58] and incorporates star structures in a technically sophisticated way. To motivate the different and non-intuitive positive dependence assumptions in Theorem 1.3, we first provide the comparison results for chain and star structures.
Regarding the notation, we write $U\uparrow_{st}V$ for random variables $U$ and $V\,,$ if $U$ is stochastically increasing (SI) in $V,$ i.e., if the conditional distribution $U|V=v$ is increasing in $v$ with respect to the stochastic order, see (13). The following result is a direct extension of the supermodular comparison of stationary discrete-time Markov processes in [58, Theorem 3.2] to a non-stationary setting.
@@
Then it follows that $X\leq_{sm}Y$ (resp. $\geq_{sm}$). In particular, $X$ and $Y$ are positive supermodular dependent.
In terms of Markov tree distributions, the underlying tree structure corresponds to a chain of nodes, see Figure 1a). Assumptions (i)-(iii) of the above proposition only refer to the bivariate distributions of $X$ and $Y$ and are therefore easy to verify. While the first two conditions are positive dependence concepts, the third condition relates to the supermodular ordering of bivariate random variables that are adjacent in the underlying chain. In contrast to higher-dimensional distributions, for bivariate distributions the supermodular order can easily be verified because, for the two-dimensional case, the supermodular order is equivalent to identical marginal distributions and the pointwise order of the associated bivariate distribution functions, see (9). Due to assumptions (i) and (ii), $X$ and $Y$ are SI in ’opposite directions’. As we discuss in Section 5, these rather odd conditions can neither be changed to the weaker notion of positive supermodular dependence nor be replaced by SI in the ’same direction’. Note that, under the assumptions of Proposition 1.1, $X$ and $Y$ are positive supermodular dependent, which is a positive dependence concept and implies, in particular, pairwise non-negative correlations.
For our main result, we make use of the following recently established supermodular ordering result which compares positive supermodular dependent random variables that are conditionally independent given a common factor variable. Such factor models may be interpreted as Markov tree distributions where the underlying tree has a star-like structure, see Figure 1b).
Lemma 1.2 (Supermodular ordering of Markovian star structures, [8, Corollary 4(i)])
Let $X=(X_{0},\ldots,X_{d})$ and $Y=(Y_{0},\ldots,Y_{d})$ be random vectors. Assume that $X_{1},\ldots,X_{d}$ are conditionally independent given $X_{0}$ and that $Y_{1},\ldots,Y_{d}$ are conditionally independent given $Y_{0}\,.$ Assume for all $i\in\{1,\ldots,d\}$ that
@@
Then it follows that $X\leq_{sm}Y\,.$ In particular, $X$ and $Y$ are positive supermodular dependent.
While for chain structures, $X$ and $Y$ must satisfy opposite SI conditions, the above result requires that $X$ and $Y$ fulfil the same SI conditions, see Section 5 and Figures 8 and 8. As our main theoretical contribution in this paper, we extend Proposition 1.1 and Lemma 1.2 to arbitrary Markov tree distributions where the underlying tree $T=(N,E)$ may have finitely or countably many nodes, see Figure 1c). To this end, let $T=(N,E)$ be a directed tree with root $0\in N\,.$ Let $L\subset N$ be the set of leaves of $T\,$, see (4). Let $P=\{\ell_{1},\ell_{2},\ell_{3}\ldots\}\subseteq N\setminus\{0\}$ with $(0,\ell_{1}),(\ell_{1},\ell_{2}),(\ell_{2},\ell_{3}),\ldots\in E$ be a path of nodes that starts with a child of the root and either terminates at a leaf node $\ell\in L$ or has infinitely many nodes. Further, let $k^{*}$ be a child of the root 00 that is not an element in the path $P$ unless it is the only child, i.e., $(0,k^{*})\in E$ with $k^{*}\notin P$ if $\operatorname{deg}(0)\geq 2$, where $\deg(n)$ denotes the degree of a node $n\in N\,,$ see Definition B.1. Then the following result holds true.
@@
Then it follows that $X\leq_{sm}Y$ (resp. $\geq_{sm}$). If additionally $(X_{0},X_{k^{*}})$ is positive supermodular dependent, then $X$ is positive supermodular dependent. Moreover, if additionally $(Y_{i},Y_{j})$ is positive supermodular dependent for $(i,j)\in E$ with $j\in P\cap L\,,$ then $Y$ is positive supermodular dependent.
[…]
Figure 2: An example that illustrates the positive dependence conditions (i) and (ii) in Theorem 1.3 for a tree $T=(N,E)$ on $12$ nodes with root 00. For $(i,j)\in E\,,$ an arrow $X_{i}\longrightarrow X_{j}$ indicates that $X_{j}$ is stochastically increasing in $X_{i}\,,$ i.e., $X_{j}\uparrow_{st}X_{i}\,.$ An arrow $X_{i}\longleftrightarrow X_{j}$ indicates that $X_{j}\uparrow_{st}X_{i}$ and $X_{i}\uparrow_{st}X_{j}\,.$ This applies similarly to the variables $(Y_{i},Y_{j})$. Note that there is no positive dependence condition between $X_{0}$ and $X_{7}$ and between $Y_{3}$ and $Y_{4}\,.$ The leaves consist of the set $L=\{2,4,5,6,10,11\}\,.$ The set $P\subseteq N$ consists of the leaf $\ell=4$ and the path $p(0,\ell)$ between the root 00 and the leaf $\ell\,,$ i.e., $P=\{1,3,4\}\,.$ The node $k^{*}$ is given by $k^{*}=7$.
The non-intuitive SI assumptions in Theorem 1.3 are illustrated in Figure 2 and can, like the supermodular comparison of bivariate distributions, be easily verified as the following remark shows. A detailed discussion of the assumptions in Section 5 proves the generality of the above theorem and establishes that none of the SI assumptions can be omitted or weakened to positive supermodular dependence. As a direct consequence of the above result, we give in Theorem 3.1 simple sufficient conditions on the bivariate copula specifications for a supermodular comparison of Markov tree distributions.
@@
    For random variable $U$ and $V,$ the SI condition $U\uparrow_{st}V$ is a positive dependence property which is equivalent to the conditional survival probability $P(U>u|V=v)$ being for all $u\in\mathbb{R}$ increasing in $v$ outside a $V$-null set that may depend on $u\,.$ In particular, this implies that $P(U>u,V>v)\geq P(U>u)P(V>v)\,,$ see Section 2.3. Many well-known families of bivariate distributions are SI, such as extreme value distributions [48], various Archimedean copulas [86], and the bivariate normal distribution for non-negative parameter of correlation [96], see also [5]. Further, the uniquely determined increasing rearrangement of a bivariate copula, recently studied in the context of dependence measures, is by construction SI, see [4, 6, 10, 107]. Since SI random vectors are invariant under increasing transformations of the components (i.e., $U\uparrow_{st}V$ implies $f(U)\uparrow_{st}g(V)$ for all increasing functions $f$ and $g$) the SI property is a copula-based dependence concept, i.e., it suffices to analyse copulas, see (5) for the notion of copula. Considering directed trees allows to incorporate asymmetric dependencies, noting that, in general, $U\uparrow_{st}V$ does not imply $V\uparrow_{st}U\,,$ see [85, Example 3.4]. Directed trees are also used for modeling causal inference, see, e.g., [31, 93].
@@
    As already mentioned, for bivariate distributions, the supermodular order can easily be verified due to its characterization by the concordance order in (9). However, for dimensions larger than $2\,,$ the supermodular order is strictly stronger than the concordance order and a verification is challenging since no small of class of functions generating the supermodular order is known. Therefore, Theorem 1.3 is meaningful because it provides a new method for constructing and comparing multivariate distributions based on bivariate building blocks, using that for bivariate distributions, various ordering result are well-known: For the bivariate normal distribution, the supermodular order corresponds to an ordering of the correlation parameter, which goes back to [106], see [7, 25, 87] for extensions to multivariate normal and elliptical distributions. A characterization of the supermodular order for bivariate Archimedean copulas in terms of their Archimedean generator and for bivariate extreme-value copulas in terms of their Pickands dependence function follows from [88, Theorem 4.4.2] and [5, Theorem 3.4], respectively. Note that it suffices to compare the underlying copulas since the supermodular order is invariant under increasing transformations.

### v3

# Comparison results for positive supermodular dependent Markov tree distributions

@@ Abstract
Positive dependencies have been compared in the literature under rather strong assumptions such as equality of conditional distributions, exchangeability, or stationarity. We establish supermodular ordering results for distributions that are Markov with respect to a tree structure. Our comparison results rely on simple stochastic monotonicity conditions and a pointwise ordering of bivariate copulas associated with the edges of the underlying tree. We also study flexibility of the marginal distributions in stochastic and convex order. As a consequence, we obtain first- and second-order stochastic dominance results for extreme order statistics and sums of positively dependent random variables. As an application, we investigate distributional robustness of the maximum of a perturbed random walk under model uncertainty. Several examples and a detailed discussion of the assumptions demonstrate the generality of our results and reveal deeper insights into non-intuitive positive dependence properties of multidimensional distributions.
Keywords bivariate copula, conditional independence, distributional robustness, factor model, hidden Markov model, graphical model, model uncertainty, pair-copula construction, perturbed random walk, positive dependence, tree-based dependence, stochastic dominance, supermodular order, vine copula model
@@ 1 Introduction
Positive dependencies play a fundamental role in both the theoretical and applied aspects of statistics. They arise naturally in a wide range of models and applications, including physics [63], economics [55], finance and risk management [29, 26], reliability theory [12], and network analysis [76, 25]. There is a large literature on concepts of positive dependence such as total positivity [32, 47, 46], conditional increasingness [59], association [34, 31], positive supermodular dependence [68], and positive quadrant dependence [51]. All these concepts involve the comparison of positively dependent random vectors with independent random vectors of the same Fréchet class.
The literature on comparing strengths of positive dependencies in a given Fréchet class includes well known parametric families of distributions, e.g. the multivariate normal distribution [57], elliptical distributions [6, 64], Archimedean copulas [60], or multivariate exponential distributions [53]. In contrast, for non-parametric classes of positively dependent distributions, ordering results are studied under quite restrictive assumptions such as equality of conditional distributions [13, 74, 75], exchangeability [70], or stationarity [38, 14]. In this paper, we establish supermodular ordering results for random vectors that are conditionally independent with respect to a tree structure. Our ordering conditions are both flexible and easy to verify, since they rely only on simple positive dependence assumptions and a pointwise ordering of the bivariate copulas associated with the edges of the underlying tree.
In many applications including reliability analysis, finance and insurance, particular important statistics are the minimum, the maximum, and the sum over the components of a random vector. The minimum may describe the first time a unit breaks down, leading to system failure. The maximum of a stochastic process is closely linked to the first passage time across a specified threshold. The aggregated components of a random vector may represent the value of a portfolio of assets or the total insurance loss. It is well known that comonotonicity describes the most extreme positive dependence for the maximum [^1] The minimum and the maximum are related by symmetry, so all results carry over to the minimum, e.g. $\min_{n}\{X_{n}\}\leq_{st}\min_{n}\{X_{n}^{c}\}.$ We will focus on the maximum in the sequel. in stochastic order and the most risky dependence structure for the sum in convex order, [^2] For real-valued random variables $S$ and $T$, the stochastic order $S\leq_{st}T$ and the convex order $S\leq_{cx}T$ are defined by $\mathbb{E}\varphi(S)\leq\mathbb{E}\varphi(T)$ for all increasing resp. convex function $\varphi$ such that the expectations exist; see Definition 2.5. i.e.,
$\max_{n}\{X_{n}\}\geq_{st}\max_{n}\{X_{n}^{c}\}\quad\text{and}\quad\sum_{n}X_{n}\leq_{cx}\sum_{n}X_{n}^{c}$
for any random vector $(X_{1},\ldots,X_{d}),$ where $X_{n}^{c}:=F_{X_{n}}^{-1}(U)$ for $U$ uniform on $(0,1);$ see [73, 21]. However, in practice, comonotonicity is too extreme to be useful, which motivates to study stochastic orderings for comparing positively dependent random variables.
In the two-dimensional case and for fixed marginals, a sufficient condition for ordering maxima and sums is the upper orthant order $\leq_{uo}$ (i.e., the pointwise order of survival functions). As a direct consequence of the ordering results in [73], $(X_{1},X_{2})\leq_{uo}(Y_{1},Y_{2})$ with $\mathsf{law}(X_{i})=\mathsf{law}(Y_{i})$ implies
$\max\{X_{1},X_{2}\}\geq_{st}\max\{Y_{1},Y_{2}\}\quad\text{and}\quad X_{1}+X_{2}\leq_{cx}Y_{1}+Y_{2}.$ (1)
These conditions are satisfied by many families of bivariate distributions and can often easily be checked. For dimension $3$ and larger, a pointwise comparison of survival functions still implies a comparison of maxima in stochastic order, but no more a comparison of aggregated components in convex order; see [58, Theorem 2.6].
A stronger dependence order that also implies a convex ordering of component sums is the supermodular order. Largest elements in supermodular order are comonotonic random vectors reflecting the worst case behavior of sums as discussed above [73]. The supermodular order is an integral stochastic order defined by comparing expectations of supermodular functions (see Definition 2.5 and Table 1). In the bivariate case, it reduces to the lower/upper orthant order, for which a verification is straightforward; see (6). In contrast, for dimension $3$ and larger, checking the supermodular order is difficult because no small class of functions generating this order is known.
Well known construction methods for random vectors that are comparable in supermodular order are based on common mixture models, majorization, sequential Markov constructions, independence, or comonotonicity [13, 14, 15, 38, 69]. The results in these references generalize the inequalities in (1) to arbitrary dimension. However, the model assumptions are quite restrictive, as they involve equality of conditional distributions or stationarity. For exchangeable random vectors, comparison results for $\max_{n}\{X_{n}\}$ in stochastic order and for the variance of $\sum_{n}X_{n}$ are derived in [70]. Note that exchangeability is closely related to conditional independence. More precisely, by de Finett’s theorem, any finite subset of an infinite sequence of exchangeable random variables is conditionally i.i.d. given a latent factor variable; see [37]. For (not necessarily conditionally i.i.d.) factor models, a supermodular ordering result was recently established in [7]. This ordering result only requires stochastic monotonicity conditions and a pointwise ordering of bivariate copulas. In particular, it extends (1) to arbitrary dimensions for conditionally independent random variables. In this paper, we generalize the supermodular ordering result for factor models to the class of Markov tree distributions which model conditional independence along tree structures; see Definition 2.2.
@@
Figure 1: Examples of Markov tree distributions: a) A Markov process in discrete time, where the underlying tree is a chain. b) Random variables $X_{1},\ldots,X_{d}$ that are conditionally independent given the common factor variable $X_{0},$ where the underlying tree has a star-like structure. c) A tree-indexed Markov process with a general underlying tree structure. d) A hidden Markov model with hidden nodes $(X_{i})_{i\in 2\mathbb{N}_{0}}$ and observable nodes $(X_{i})_{i\in 2\mathbb{N}_{0}+1}$. Each model is uniquely determined by the univariate marginal distributions specifying the nodes and by the bivariate copulas $(B_{ij})_{(i,j)\in E}$ specifying the edges of the underlying tree $T=(N,E);$ see Proposition 2.4.
Markov tree distributions are a flexible and interpretable class of models because they admit a simple representation in terms of bivariate distributions. More precisely, for a directed tree $T=(N,E)$ of nodes $N$ and edges $E\subset N\times N,$ a fundamental property of Markov tree distributions is that they can be constructed by a sequence $F=(F_{n})_{n\in N}$ of univariate distribution functions and a sequence $B=(B_{e})_{e\in E}$ of bivariate copulas. While $F_{n}$ determines the distribution of $X_{n},$ the copula $B_{e},$ $e=(i,j)\in E,$ describes the dependence structure between the random variables $X_{i}$ and $X_{j}$ that are adjacent in $T$; see Figure 1 and Proposition 2.4. Hence, Markov tree distributions are pair-copula constructions—and analyzing their distributional properties reduces to studying their specifications. We write $\mathcal{M}(F,T,B)$ for the Markov tree distribution with these specifications. Note that Markov tree distributions are the cornerstone of vine copulas, which have become very popular in the last decade [16, 23, 24]. More precisely, if Lebesgue densities exist, Markov tree distribution are vine copula models with only one layer.
For continuous marginal distribution functions, it is clear that Markov-tree distributed random variables $(X_{n})_{n\in N}\sim\mathcal{M}(F,T,B)$ are independent if and only if all the bivariate copulas $B_{e}$ coincide with the independence copula. Similarly, $(X_{n})_{n\in N}$ are comonotonic if and only if all $B_{e}$ coincide with the comonotonicity copula. The focus of this paper is to derive simple conditions on the specifications $(F_{n})_{n\in N}$ and $(B_{e})_{e\in E}$ to model and order positive dependencies between the extreme cases of independence and comonotonicity.
We contribute to the literature on positive dependence orderings as follows.
  1. 1.
    Our main result, Theorem 1.3, compares Markov tree distributions in supermodular order. In particular, it extends (1) to conditionally independent models along tree structures noting that $(X_{n})\leq_{sm}(Y_{n})$ implies [^3] We generally assume integrable random variables so that a comparison of the aggregated components in convex order makes sense.
    $\max_{n}\{X_{n}\}\geq_{st}\max_{n}\{Y_{n}\}\quad\text{and}\quad\sum_{n}X_{n}\leq_{cx}\sum_{n}Y_{n}.$
    The non-intuitive positive dependence assumptions and the ordering conditions in Theorem 1.3 can easily be checked because they rely only on the bivariate distributions related to the edges of the underlying tree; see Remark 1.5 (a) and (b). In Corollary 1.4, we formulate a version of our main result based on the bivariate copula specifications.
  2. 2.
    In Theorem 3.1, we additionally allow for flexibility of the marginal distributions in stochastic order and establish a comparison of Markov tree distributions with respect to the increasing supermodular order. This ordering continues to imply first- and second-order stochastic dominance results for extreme value statistics and sums of random variables; see (5). Similarly, Theorem 3.5 permits flexibility of the marginal distributions in convex order, leading to a comparison result in the directionally convex order. The latter order implies a comparison of the aggregated components in convex order.
  3. 3.
    Theorem 1.3 is general in the sense that none of the positive dependence conditions can be omitted or weakened to positive supermodular dependence; see Proposition A.2. Hence, our results reveal new insights into ordering properties of multidimensional distributions. Theorem 1.3 also allows us to compare some negative dependencies, as not all bivariate dependence specifications are required to satisfy stochastic monotonicity conditions. In Section 4, we investigate a particular property of star structures, for which arbitrary dependence specifications can be compared with positive dependencies in supermodular order; see Proposition 4.2. Interestingly, the underlying ordering conditions are consistent with dependence measures such as Chatterjee’s rank correlation [8, 11]. As we show, this ordering result for star structures does not extend to chains and, consequently, cannot be generalized to Markov tree distributions.
  4. 4.
    Our contributions admit various applications to Markovian structures, including hidden Markov models, as we study in Section 5. Corollary 5.1 is a version of our comparison results in the context of hidden Markov models, where the observations are modeled as conditionally independent given the latent process. As an illustrative example, we consider the maximum of the observations and study its distributional robustness in stochastic order; see Theorem 5.4. For simplicity, the hidden process is modeled as a random walk—a setting commonly referred to as a perturbed random walk. While the existing literature on maxima of perturbed random walks focuses on their asymptotic behavior [10, 39, 40], our robustness results pertain to a finite number of observations.
1.1 Main results
@@
To motivate the non-intuitive positive dependence assumptions in Theorem 1.3 below, we first formulate comparison results for chain and star structures, both of which are based on asymmetric positive dependence conditions.
Regarding the notation, we write $V\uparrow_{st}U$ for random variables $U$ and $V,$ if $V$ is stochastically increasing (SI) in $U;$ see (7). The supermodular order $X\leq_{sm}Y$ is defined by $\mathbb{E}f(X)\leq\mathbb{E}f(Y)$ for all $b$-bounded, supermodular functions $f$; see Definition 2.5, where a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to be supermodular if $f(x)+f(y)\leq f(x\wedge y)+f(x\vee y)$ for all $x,y\in\mathbb{R}^{n}.$ Here, $\wedge$ and $\vee$ denote the componentwise minimum and maximum, respectively. Further, a random vector is said to be positive supermodular dependent (PSMD) if it exceeds an independent random vector in supermodular order; see Definition 2.6. The following result is a direct extension of [38, Theorem 3.2] from stationary discrete-time Markov processes to a non-stationary setting. In the language of Markov tree distributions, the underlying tree structure corresponds to a chain of nodes; see Figure 1a).
@@
Then it follows that $X\leq_{sm}Y$ (resp. $\geq_{sm}$). In particular, $X$ and $Y$ are PSMD.
Assumptions (i) – (iii) in the above proposition refer only to bivariate distributions and are therefore easy to verify; see Remark 1.5 (a) and (b). Due to the first two conditions, $X$ and $Y$ are SI in ’opposite directions’. As we discuss in the Proposition A.2 and Example 6.1, these rather non-intuitive positive dependence assumptions cannot be weakened to PSMD, nor can they be replaced with SI in the ’same direction’. Note that, under the assumptions of Proposition 1.1, the entire processes $X$ and $Y$ are PSMD.
For the proof of our main result, we use the following recently established ordering result for factor models where conditionally independent random variables are compared with respect to the supermodular order; see [7, Corollary 4(i)]. Such factor models can be interpreted as Markov tree distributions where the underlying tree has a star-like structure; see Figure 1b).
Lemma 1.2 (Supermodular ordering of Markovian star structures)
Let $X=(X_{0},\ldots,X_{d})$ and $Y=(Y_{0},\ldots,Y_{d})$ be random vectors. Assume that $X_{1},\ldots,X_{d}$ are conditionally independent given $X_{0}$, and that $Y_{1},\ldots,Y_{d}$ are conditionally independent given $Y_{0}.$ Assume for all $i\in\{1,\ldots,d\}$ that
@@
Then it follows that $X\leq_{sm}Y.$ In particular, $X$ and $Y$ are PSMD.
While for chain structures, $X$ and $Y$ are required to satisfy opposite SI conditions, it is important to emphasize that the above result demands $X$ and $Y$ to fulfill identical SI conditions; see Example 6.2.
To extend Proposition 1.1 and Lemma 1.2 to Markov tree distributions, it becomes necessary to combine the SI conditions in a suitable way. To this end, let $T=(N,E)$ be a directed tree with root $0\in N$, which may have finitely or countably many nodes; see Figure 1c). Denote by $L\subset N$ the set of leaves of $T.$ Let $P=\{\ell_{1},\ell_{2},\ell_{3}\ldots\}\subseteq N\setminus\{0\}$ with $(0,\ell_{1}),(\ell_{1},\ell_{2}),(\ell_{2},\ell_{3}),\ldots\in E$ be a path of nodes that starts with a child of the root and either terminates at a leaf node $\ell\in L$ or has infinitely many nodes. Further, let $k^{*}$ be a child of the root 0 that does not belong to the path $P$, unless it is the only child, i.e., $(0,k^{*})\in E$ with $k^{*}\notin P$ whenever $\operatorname{deg}(0)\geq 2$. We refer to Section 2.1.1 and Definition B.1 for a formal specification of the notation and the terms. The following theorem is our main result.
@@
Then it follows that $X\leq_{sm}Y$ (resp. $\geq_{sm}$). If additionally $(X_{0},X_{k^{*}})$ is PSMD, then $X$ is PSMD. Moreover, if additionally $(Y_{i},Y_{j})$ is PSMD for $(i,j)\in E$ with $j\in P\cap L,$ then $Y$ is PSMD.
The non-intuitive SI conditions in Theorem 1.3 are illustrated in Figure 2. A detailed discussion of the assumptions provided in Appendix A proves the generality of the above theorem and establishes that none of the SI conditions can be omitted or relaxed to PSMD. The technical proof of Theorem 1.3, presented in Appendix B.1, builds on the proof for ordering stationary Markov processes in [38] and extends it by incorporating star structures. This requires a series of auxiliary results and additional notation to navigate appropriately through the tree. Note that Theorem 1.3 compares dependence structures of Markov tree distributions for arbitrary univariate marginals. Related results for Poisson and Binomial random vectors are given in [22, 48].
As a consequence of the above result, the following corollary provides simple sufficient conditions on the bivariate copula specifications for a supermodular comparison of Markov tree distributions. We denote by $\leq_{lo}$ the lower orthant (i.e., pointwise) order of distribution functions. Further, a bivariate distribution function is said to be SI (CI) if there exists a bivariate random vector $(U,V)$ having this distribution function such that $V\uparrow_{st}U$ (and $U\uparrow_{st}V$); see Definition 2.6.
Corollary 1.4 (Supermodular ordering based on bivariate tree specifications)
Let $X\sim\mathcal{M}(F,T,B)$ and $Y\sim\mathcal{M}(F,T,C).$ Assume for all $e\in E$ that
  1. (i)
    $B_{e}$ is SI,
  2. (ii)
    $C_{e}$ is CI,
  3. (iii)
    $B_{e}\leq_{lo}C_{e}$ (resp. $\geq_{lo}$).
Then it follows that $X\leq_{sm}Y$ (resp. $\geq_{sm}$). In particular, $X$ and $Y$ are PSMD.
$X_{0}$$X_{1}$$X_{2}$$X_{3}$$X_{4}$$X_{5}$$X_{6}$$X_{7}$$X_{8}$$X_{9}$$X_{10}$$X_{11}$
$Y_{0}$$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$$Y_{5}$$Y_{6}$$Y_{7}$$Y_{8}$$Y_{9}$$Y_{10}$$Y_{11}$
Figure 2: An example that illustrates the positive dependence conditions (i) and (ii) in Theorem 1.3 for a tree $T=(N,E)$ on 1212 nodes with root 0. An arrow $U\longrightarrow V$ indicates $V\uparrow_{st}U;$ similarly, an arrow $U\longleftrightarrow V$ indicates $V\uparrow_{st}U$ and $U\uparrow_{st}V.$ Note that there is no positive dependence condition between $X_{0}$ and $X_{7}$ and between $Y_{3}$ and $Y_{4}.$ The leaves consist of the set $L=\{2,4,5,6,8,10,11\}.$ The set $P\subseteq N$ consists of the leaf $\ell=4$ and the path $p(0,\ell)$ between the root 0 and the leaf $\ell,$ i.e., $P=\{1,3,4\}.$ The root’s child $k^{*}$ is given by $k^{*}=7$.
@@
    For random variables $U$ and $V,$ the SI condition $V\uparrow_{st}U$ is a positive dependence concept, satisfied by many well known bivariate distributions. For example, extreme value distributions [35], various Archimedean copulas [60], and the bivariate normal distribution for non-negative correlation [66] are SI; see [4] for an overview. Further, the uniquely determined increasing rearrangement of a bivariate copula, recently studied in the context of dependence measures, is by construction SI; see [5, 72]. Since SI random vectors are invariant under increasing transformations of the components (i.e., $U\uparrow_{st}V$ implies $f(U)\uparrow_{st}g(V)$ for all increasing functions $f$ and $g$; see [17, Proposition 2.12]) the SI property is a copula-based dependence concept. Hence, it suffices to analyze copulas; see (3) for the notion of copula.
@@
    For bivariate distributions, the supermodular order can be verified by a pointwise comparison of the underlying copulas. This follows from (6) and from the invariance of the supermodular order under increasing transformations [69, Theorem 9.A.9(a)]. In contrast, for dimensions greater than $2,$ the supermodular order is strictly stronger than the orthant orders, making a verification challenging since no small class of functions generating the supermodular order is known [61]. This is why Theorem 1.3 is particularly meaningful: it provides a novel method for constructing and comparing PSMD multivariate distributions based on bivariate building blocks, leveraging well known ordering results for bivariate distributions. For instance, in the case of the bivariate normal distribution, the supermodular order corresponds to an ordering of the correlation parameter [61, Theorem 3.13.5]; see [6] for an extension to elliptical distributions. For bivariate Archimedean copulas, a characterization of the supermodular order follows from [62, Theorem 4.4.2]. For bivariate extreme value copulas, the supermodular order is characterized in [4, Theorem 3.4].